OpenStudy (anonymous):

Please help! Which function does not have a vertical asymptote? A. y=5x/1-x^2 B. y=x/1+x^2 C. y=5x-1/x D. y=5x/x+x^2

OpenStudy (solomonzelman):

Do you know what is a vertical asymptote is?

OpenStudy (solomonzelman):

HINT: a vertical asymptote occurs when x cannot equal to something.

OpenStudy (solomonzelman):

@ali1029, so what do you think?

OpenStudy (anonymous):

i am so sorry @SolomonZelman, I had to help my little brother with something. I got your message but can you please explain to me how you got that answer?

OpenStudy (solomonzelman):

I gave you a hint above....

OpenStudy (anonymous):

well yes, i saw that but what steps did you take to find the answer? did you just put a 0 in for all of the x's?

OpenStudy (anonymous):

it says that the answer is B.

OpenStudy (solomonzelman):

Yup! and in choices x is either 1 or 0, but in a x could be anything.

terenzreignz (terenzreignz):

... (_^^) All these choices have denominators which are functions of x, right?

OpenStudy (solomonzelman):

I mean a is the answer.

OpenStudy (solomonzelman):

And you already said y.

terenzreignz (terenzreignz):

The answer is B though :)

OpenStudy (solomonzelman):

I know it is. I really though it's A, sorry!

terenzreignz (terenzreignz):

Look @ali1029 all these choices have their denominators, aye? So... there's something special about the denominator in choice B... do you know what it is?

OpenStudy (anonymous):

it's okay @SolomonZelman . & no, I'm not too sure what makes the denominator special :o @terenzreignz

terenzreignz (terenzreignz):

It's simple ^_^ The denominator in B could NEVER be equal to zero, no matter what value x takes. 1+x^2 1 is positive, and x^2 is always positive (or zero) either way, the denominator in choice B is always going to be at least 1. Does that make sense?

terenzreignz (terenzreignz):

Look at choice A. 1 - x^2 it becomes zero when x = 1 or x = -1 Look at choice C x clearly, this is zero when, well, x = 0 LOL Look at choice D x+x^2 This is zero when x = 0 or x = -1 but choice B, no matter what value x takes, 1+x^2 is never zero, and a vertical asymptote is something that results from a zero-denominator ^_^

OpenStudy (anonymous):

Ohh, i see, that completely makes sense now! thank you so much :) if you dont mind, do you think you could help me with a few more problems i have on this topic? i really need to make a good grade :o

terenzreignz (terenzreignz):

Maybe one more. I'm sleepy ^_^

OpenStudy (anonymous):

oh haha i just woke up :p

OpenStudy (anonymous):

well, what would the horizontal asymptote be for the function: y=x^2/x^2-9?

terenzreignz (terenzreignz):

oh, lol, is this a calculus question?

OpenStudy (anonymous):

i think so, im in pre calculus :o

terenzreignz (terenzreignz):

Can we use limits?

OpenStudy (anonymous):

i'm not sure, i dont really know what those are lol, heres the official question

terenzreignz (terenzreignz):

Well, I'll give you a shortcut then, but this only works if the highest powers in both the numerator and denominator are equal, ok? In this case, in the numerator, the highest power is 2, same in the denominator, right?

OpenStudy (anonymous):

correct!

terenzreignz (terenzreignz):

Then, the asymptote at infinity would simply be the quotient of their coefficients, both of which are 1, s....?

OpenStudy (anonymous):

so the horizontal asymptote would just be 1?

terenzreignz (terenzreignz):

Yup.

terenzreignz (terenzreignz):

Well... I'm signing off now... Good luck with your grade :) -------------------------------- Terence out

OpenStudy (anonymous):

Thank you so much, you're amazing! :)

OpenStudy (solomonzelman):

he taught me the limits, or almost did....

OpenStudy (anonymous):

haha yep :p